is considered circular with a diameter of 2b. A participant starts his vehicle manages to increases his speed at a constant rate of α. Obtain a simplified expression between his velocity and his acceleration as a function of time


1
Expert's answer
2021-05-25T09:29:58-0400

We know that circuler path particle move

Then two force are work

"F_{ext}=\frac{mv^2}{b}"

"F{ext}=m\alpha"

"m\alpha=\frac{mv^2}{b}"

"\alpha=\frac{v^2}{b}"

Where b radius of circuler path